OpenStudy (anonymous):

3x-5y+z=8, 4y-z=10 ,7x+y=4

OpenStudy (texaschic101):

I would solve by elimination....are you familiar with that method ?

OpenStudy (anonymous):

x = 11/5, y = -57/5, z = -278/5

OpenStudy (anonymous):

can you guys solve step by step

OpenStudy (texaschic101):

wait...you can't use elimination because the 2nd equation has no x value...you have to use substitution

OpenStudy (anonymous):

come on help me this out please

OpenStudy (abb0t):

you could use cramers rule - matrices.

OpenStudy (anonymous):

whats that rule omg.. never heard about that

OpenStudy (anonymous):

Solve the following system: {3 x-5 y+z = 8 | (equation 1) 4 y-z = 10 | (equation 2) 7 x+y = 4 | (equation 3) Swap equation 1 with equation 3: {7 x+y+0 z = 4 | (equation 1) 0 x+4 y-z = 10 | (equation 2) 3 x-5 y+z = 8 | (equation 3) Subtract 3/7 × (equation 1) from equation 3: {7 x+y+0 z = 4 | (equation 1) 0 x+4 y-z = 10 | (equation 2) 0 x-(38 y)/7+z = 44/7 | (equation 3) Multiply equation 3 by 7: {7 x+y+0 z = 4 | (equation 1) 0 x+4 y-z = 10 | (equation 2) 0 x-38 y+7 z = 44 | (equation 3) Swap equation 2 with equation 3: {7 x+y+0 z = 4 | (equation 1) 0 x-38 y+7 z = 44 | (equation 2) 0 x+4 y-z = 10 | (equation 3) Add 2/19 × (equation 2) to equation 3: {7 x+y+0 z = 4 | (equation 1) 0 x-38 y+7 z = 44 | (equation 2) 0 x+0 y-(5 z)/19 = 278/19 | (equation 3) Multiply equation 3 by 19: {7 x+y+0 z = 4 | (equation 1) 0 x-38 y+7 z = 44 | (equation 2) 0 x+0 y-5 z = 278 | (equation 3) Divide equation 3 by -5: {7 x+y+0 z = 4 | (equation 1) 0 x-38 y+7 z = 44 | (equation 2) 0 x+0 y+z = -278/5 | (equation 3) Subtract 7 × (equation 3) from equation 2: {7 x+y+0 z = 4 | (equation 1) 0 x-38 y+0 z = 2166/5 | (equation 2) 0 x+0 y+z = -278/5 | (equation 3) Divide equation 2 by -38: {7 x+y+0 z = 4 | (equation 1) 0 x+y+0 z = -57/5 | (equation 2) 0 x+0 y+z = -278/5 | (equation 3) Subtract equation 2 from equation 1: {7 x+0 y+0 z = 77/5 | (equation 1) 0 x+y+0 z = -57/5 | (equation 2) 0 x+0 y+z = -278/5 | (equation 3) Divide equation 1 by 7: {x+0 y+0 z = 11/5 | (equation 1) 0 x+y+0 z = -57/5 | (equation 2) 0 x+0 y+z = -278/5 | (equation 3) Collect results: Answer: | | {x = 11/5 y = -57/5 z = -278/5 That is elimination.

OpenStudy (anonymous):

wow so nice and useless

OpenStudy (abb0t):

spam.

OpenStudy (texaschic101):

y = -7x + 4 now sub -7x + 4 in for y in the 2nd equation 4y - z = 10 4(-7x + 4) - z = 10 -28x + 16 - z = 10 -28x - z = 10 - 16 -28x - z = -6 -z = 28x - 6 z = -28x + 6 now sub -7x + 4 in for y and -28x + 6 in for zin the 2st equation 3x - 5y + z = 8 3x - 5(-7x + 4) -28x + 6 = 8 3x + 35x - 20 - 28x + 6 = 8 38x - 28x - 14 = 8 10x = 8 + 14 10x = 22 x = 22/10 = 11/5 now sub 11/5 in for x in the 3rd equation 7x + y = 4 7(11/5) + y = 4 77/5 + y = 4 y = 4 - 77/5 y = 20/5 - 77/5 y = - 57/5 now sub -57/5 in for y in the 2nd equation 4y - z = 10 4(-57/5) - z = 10 -228/5 - z = 10 -z = 10 + 228/5 -z = 50/5 + 228/5 -z = 278/5 z = -278/5 check the answers 3x - 5y + z = 8 3(11/5) - 5(-57/5) - 278/5 = 8 33/5 + 285/5 - 278/5 = 8 318/5 - 278/5 = 8 40/5 = 8 8 = 8 (correct) jprulz is correct......x = 11/5, y = -57/5, z = -278/5

OpenStudy (anonymous):

wow nice work bros love you guys all

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

thank you so much texaschic XP